# The galaxy-IGM connection in thesan: observability and information content of the galaxy-Lyman- cross-correlation at z6

### v1

@@ 5.4 Measuring the GaLCC for Helium reionization
Finally, we address one last question concerning the GaLCC, namely the possibility to extend the use of this probe to the Helium epoch of reionization (HeEoR), replacing the HāI Ly line with the HeāII one and galaxies with QSOs. There are a number of factors, both contributing and hindering this endeavor, to take into account. For this reason, a full examination of this problem requires a dedicated work, that we plan to deliver in the future. However, we briefly discuss here the main considerations and insights from our results.
@@
The HeEoR occurs at much lower redshift than the hydrogen reionization discussed so far (approximately at , see e.g.Ā Worseck etĀ al., 2016; Davies etĀ al., 2017; Worseck etĀ al., 2019; Makan etĀ al., 2021, 2022), where the number density of QSO is higher and, therefore, there are potentially more sightlines that can be used to build the cross-correlation. For example, WEAVE 444The WHT Enhanced Area Velocity Explorer, https://ingconfluence.ing.iac.es/confluence//display/WEAV will soon more than double the number of sightlines observed at (covering several thousand square degrees). Although the wavelength range of WEAVE does not cover the redshifted HeāII Ly wavelength of Ć Ć (for ), it can be used to identify sightlines clean from contamination of higher- HāI Lyman continnum for follow up. At the same time, eBOSS555https://www.sdss4.org/surveys/eboss/, 4MOST666https://www.eso.org/sci/facilities/develop/instruments/4MOST.html and EUCLID777https://sci.esa.int/web/euclid/ are also expected to significantly increase the number of known QSOs (although, again, without the necessary wavelength coverage). However, observations of the HeāII Ly forest are significantly more challenging than those of the HāI counterpart, due to the aforementioned contamination and to the necessity of spaced-based observations (since the atmosphere is opaque at the FUV wavelengths where the helium forest is found). Additionally, the rarity of the sources driving the HeEoR implies that a much larger area around each sightline needs to be covered in order to identify sources contributing to the GaLCC (since they produce much larger ionised bubbles, e.g.Ā Compostella etĀ al., 2013). The situation is worsened by the fact that QSO emission is thought to be strongly anisotropic, which introduces geometrical (unobservable) weights to the contribution of each source around the line of sight. However, QSOs are also much brighter than galaxies, therefore imprinting a much stronger signature on the surrounding Ly forest which might make up for the larger sensitivity to source-to-source variations. Finally, QSOs reside in the largest overdensities (e.g.Ā Efstathiou & Rees, 1988; Volonteri & Rees, 2006; Costa etĀ al., 2014; Costa, 2024), therefore boosting the HeāIII recombination rate and suppressing the HeāII Ly transmissivity. Overall, this appears a complex problem deserving a dedicated study in the future.
@@
While the source-Ly cross-correlation has never been studied in the HeEoR, there have been somewhat similar attempts. For instance, Schmidt etĀ al. (2017) presented a survey of QSOs at designed specifically to detect the HeāII transverse proximity effect, i.e.Ā the enhancement in the nearby Ly transmissivity due to the quasar ionizing photon production, that resulted in 66 āscience-gradeā spectra probing 4 different foreground QSOs. Their goal was to constrain the lifetime and obscuration of individual QSOs, and therefore analyzed individual features in the spectra (Schmidt etĀ al., 2018, finding evidence of such proximity effect). However, a similar, expanded sample would be a promising tool to measure the source-Ly cross-correlation and put new constraints on the HeEoR, pushing forward one of the least studied periods in the history of structure formation. This acquired even more importance in recent times, since it was pointed out by Basu etĀ al. (2024) that our current model of QSO luminosity function evolution yields results that are somewhat in tension with the (still sparse) constraints on the HeEoR.
@@ 5.5 Implications for the relation between sightline effective optical depth and galaxy density
The GaLCC is closely related to the relationship between the total Ly effective optical depth of a sightline and the number density of galaxies around it, that has been recently measured in quasar sightlines (Becker etĀ al., 2018; Kashino etĀ al., 2020; Christenson etĀ al., 2021, 2023; Ishimoto etĀ al., 2022), including some of the most transparent and most opaque LOS known at ā. We investigate the implications of the GaLCC and the predictions of thesan regarding this quantity in a companion paper (Garaldi & Bellscheidt, in preparation).
@@ 6 Summary and Conclusions
The galaxy-Ly cross-correlation (GaLCC) has been proposed as a novel way to constrain the properties and timing of cosmic reionization. In this paper we have (i) thoroughly investigated the extent to which a different observational limitations prevent us from extracting the intrinsic GaLCC signal, as well as (ii) further clarified the physical origin of the signal and its dependence on galaxy properties and their environment. In order to do so, we have employed the thesan suite of simulations (specifically, the thesan-1 flagship run) to produce synthetic Ly forest lines of sight. Our main results are the following:
ā¢
@@
ā¢
We predict the recent observations by Kashino etĀ al. (2023) to be dominated by stochasticity, and therefore we caution against interpreting the in a quantitative or qualitative way. However, the full EIGER survey is poised to faithfully recover the intrinsic GaLCC signal if the same redshift bins are used, therefore covering a total length of .
ā¢
The galaxy selection method appears to be largely irrelevant for the GaLCC. We do not find differences when the cross-correlation is computed using the galaxies with the largest stellar mass, SFR, gas metallicity (as a proxy for their CāIV absorption) and OāIII flux. This stems from the fact that thesan predicts galaxies at to lie on the galaxy main sequence and the mass-metallicity relation, and the OāIII flux to be correlated with the galaxy SFR. Hence, galaxies selected to be extreme in one of these quantities are typically extreme also in the others.
ā¢
@@
ā¢
The excess transmitted flux at intermediate scales is driven by galaxy overdensities rather than individual bright objects. Therefore, the GaLCC does not trace ionised bubbles produced by individual galaxies, but rather large-scales collective bubbles. This clarifies why Garaldi etĀ al. (2022) found that, simultaneously, (i) even when ionizing photons are allowed to escape only from small () galaxies, the GaLCC signal is not affected, and (ii) the flux excess is larger around massive (, ) galaxies (i.e.Ā when only such galaxies are used to compute the GaLCC). The emerging picture is the following: small galaxies provide the majority of the ionizing photons required to induce an excess Ly transmission, but individual sources are not able to produce an observable signature because of the large āstructure formation noiseā. Therefore, only large groups of such galaxies can produce an observable signature. These groups typically reside in overdense regions, as do highly biased sources like massive galaxies. Hence, using these massive objects to compute the GaLCC effectively selects conglomerates of small galaxies, producing a stronger signal. At the same time, if larger galaxies are not allowed to emit ionizing photons, this does not significantly affect the GaLCC, since the majority of ionising photons is produced by the numerous, nearby smaller galaxies.
ā¢
The most opaque (transparent) sightlines to Ly photons show larger and closer to the galaxy (smaller and farther from the galaxy) peaks in the GaLCC. This is due to: (i) the fact that the transmitted flux is normalized by its average in the sightline, which is larger in transparent sightlines; and (ii) tho overall lower ionising photons density in opaque sightlines renders more difficult to reach the high ionization levels required to allow Ly transmission, and therefore moves the peak transmissivity at smaller distances from the galaxy.
ā¢
@@
ā¢
The redshift evolution of the GaLCC peak position is compatible with the evolving density field in a fully-ionised IGM, as expected around the brightest sources of reionization. Quantitatively, the position is approximately times smaller than the mean free path evolution extrapolated from Worseck etĀ al. (2014, which is computed from measurements in the post-reionization IGM).
ā¢
@@
Overall, our analysis sheds new light on the galaxy-Ly cross correlation. It clarifies further the physical origin and the thoroughly investigate the observational requirements for a faithful determination of this signal, reviewing the ability of current and forthcoming surveys to deliver a faithful measurement. We also provide an extensive discussion of future possible extension of this type of studies. As such, this paper represents a step further to a deeper understanding of the galaxy-IGM interplay during the first billion years of the Universe, a step direly needed to shed light on the many questions being on a daily basis by observations of the reionizing Universe.
@@ Acknowledgements
EG wishes to express his gratitude towards Koki Kakiichi and Christopher Cain for discussions that inspired part of this work, as well as towards Louise Welsh, Daichi Kashino, Valentina DāOdorico and Matthew Pieri for useful discussions. We are thankful to the community developing and maintaining software packages extensively used in our work, namely: matplotlib (Hunter, 2007), numpy (Walt etĀ al., 2011), scipy (Jones etĀ al., 2001), cmasher (van der Velden, 2020) and CoReCon (Garaldi, 2023).
@@
All simulation data, including snapshots, group and subhalo catalogues, merger trees, and high time cadence Cartesian outputs are publicly available at www.thesan-project.com and thoroughly described in Garaldi etĀ al. (2024), including the additional synthetic spectra developed for this work.
@@
We list here the authors contribution following the CRediT888https://www.elsevier.com/researcher/author/policies-and-guidelines/credit-author-statement system. EG: conceptualization, methodology, software, formal analysis, validation, writing ā original draft, writing ā review and editing, visualization, supervision, project administration. VB: software, formal analysis, writing ā review and editing.
@@ Appendix A Public release of 300 additional sightlines and lightcone-like LOS
Redshift | - | Redshift of each segment. Notice that this is the redshift the segment would have if the LOS was extracted from a real lightcone (i.e.Ā without piecewise-constant approximation). To know the actual redshift, see the field āSnapshotā |
@@
The 300 additional sightlines have the same data format as the original 150 LOS described in Section 3.10 of Garaldi etĀ al. (2024). They are located at Thesan/Thesan-1/postprocessing/los_zdir. In this directory, there is one file per snapshot containing all LOS. These files are named rays_zdir_NNN.hdf5, where NNN is the 0-left-padded snapshot number (starting from 54).
@@
The lightcone-like LOS are all contained in a single file located at Thesan/Thesan-1/postprocessing/los_lc/rays_lc.hdf5. They share the same format as the fixed-redshift LOS, but contain some additional datasets, reported in TableĀ 2.
Appendix B Impact of forcing the flux to be positive
@@
In our investigation of the impact of noise on the GaLCC in Sec.Ā 3.3 we do not impose the flux in each pixel to be positive after the addition of noise. This choice stems from the consideration that errors in the quasar continuum estimation can lead to negative flux values. Here we investigate the impact of this choice. Interestingly, we find that forcing the flux to be positive makes the cross-correlation signal much less resilient to increasing noise in the spectrum, simultaneously changing the way it is affected by the noise. We show the impact of forcing the flux in each pixel to be non-negative (essentially by setting all negative flux values to zero) on the cross-correlation in Fig.Ā 21. The difference with Fig.Ā 7 is striking. In this case, the noise does not simply broaden the range of recovered cross-correlation values around the intrinsic one, but instead suppresses the variations (both at small and intermediate scales) and eventually flattens the signal. Therefore, even in the case of moderate S/N (approximately 30%), the recovered signal is far from the intrinsic one. Even more worryingly, the recovered signal appears similar to the intrinsic one at a later time (or, equivalently, larger ionised fraction), potentially preventing the ability to use this probe to pinpoint the completion of the reionization process. The reason of this behaviour change lies in the fact that imposing a positive flux introduces an asymmetry, since now noise can only increase the total flux, unlike in the previous case. Therefore, the noise does not average out when computing the cross-correlation. Instead, it effectively increases the flux regardless of the position of nearby galaxies, hence flattening the cross-correlation.
Appendix C Comparison of effective, median and middle redshift
@@
In Sec.Ā 3.6 we have computed an effective redshift of the GaLCC computed over a broad redshift window by identifying the closest GaLCC computed at fixed redshift (or, equivalently, computed over a narrow redshift range). In Fig.Ā 22 we show how such effective redshift tracks the central redshift of the window used (, top panel) and the median redshift of all galaxies used (, bottom panel), for all combinations of and (the latter shown by the color of the points). In the bottom panel, points are much closer to the one-to-one line (dashed grey line), showing that tracks much better the effective redshift of the cross-correlation computed from the lightcone-like los.
@@
Interestingly, appears to almost-fully capture the impact of large redshift windows, as can be seen in the bottom panel of Fig.Ā 22, where all light blue/green points lie close to the one to one line (grey dashed line). This is not true for . At , the scatter around the line increases, indicating that the effective redshift is a worse predictor of than at lower redshift. This is due to the increasing noise in the cross-correlation that makes the identification of an effective redshift more dependent on the noise (arising from the lower number of galaxies fulfilling our selection criteria (i.e.Ā ) decreases at earlier and earlier times. We note that this would be true for any selection criteria based on galaxy mass or brightness.
Appendix D GaLCC evolution fit

### v2

@@ 5.4 Measuring the GaLCC for Helium reionization
Finally, we address one last question concerning the GaLCC, namely the possibility to extend the use of this probe to the Helium epoch of reionization (HeEoR), replacing the H I Ly line with the He II one and galaxies with QSOs. There are a number of factors, both contributing and hindering this endeavor, to take into account. For this reason, a full examination of this problem requires a dedicated work, that we plan to deliver in the future. However, we briefly discuss here the main considerations and insights from our results.
@@
The HeEoR occurs at much lower redshift than the hydrogen reionization discussed so far (approximately at , see e.g. Worseck et al., 2016; Davies et al., 2017; Worseck et al., 2019; Makan et al., 2021, 2022), where the number density of QSO is higher and, therefore, there are potentially more sightlines that can be used to build the cross-correlation. For example, WEAVE 555The WHT Enhanced Area Velocity Explorer, https://ingconfluence.ing.iac.es/confluence//display/WEAV will soon more than double the number of sightlines observed at (covering several thousand square degrees). Although the wavelength range of WEAVE does not cover the redshifted He II Ly wavelength of Å Å (for ), it can be used to identify sightlines clean from contamination of higher- H I Lyman continnum for follow up. At the same time, eBOSS666https://www.sdss4.org/surveys/eboss/, 4MOST777https://www.eso.org/sci/facilities/develop/instruments/4MOST.html and EUCLID888https://sci.esa.int/web/euclid/ are also expected to significantly increase the number of known QSOs (although, again, without the necessary wavelength coverage). However, observations of the He II Ly forest are significantly more challenging than those of the H I counterpart, due to the aforementioned contamination and to the necessity of spaced-based observations (since the atmosphere is opaque at the FUV wavelengths where the helium forest is found). Additionally, the rarity of the sources driving the HeEoR implies that a much larger area around each sightline needs to be covered in order to identify sources contributing to the GaLCC (since they produce much larger ionized bubbles, e.g. Compostella et al., 2013). The situation is worsened by the fact that QSO emission is thought to be strongly anisotropic, which introduces geometrical (unobservable) weights to the contribution of each source around the line of sight. However, QSOs are also much brighter than galaxies, therefore imprinting a much stronger signature on the surrounding Ly forest which might make up for the larger sensitivity to source-to-source variations. Finally, QSOs reside in the largest overdensities (e.g. Efstathiou & Rees, 1988; Volonteri & Rees, 2006; Costa et al., 2014; Costa, 2024), therefore boosting the He III recombination rate and suppressing the He II Ly transmissivity. Overall, this appears a complex problem deserving a dedicated study in the future.
@@
While the source-Ly cross-correlation has never been studied in the HeEoR, there have been somewhat similar attempts. For instance, Schmidt et al. (2017) presented a survey of QSOs at designed specifically to detect the He II transverse proximity effect, i.e. the enhancement in the nearby Ly transmissivity due to the quasar ionizing photon production, that resulted in 66 ‘science-grade’ spectra probing 4 different foreground QSOs. Their goal was to constrain the lifetime and obscuration of individual QSOs, and therefore analyzed individual features in the spectra (Schmidt et al., 2018, finding evidence of such proximity effect). However, a similar, expanded sample would be a promising tool to measure the source-Ly cross-correlation and put new constraints on the HeEoR, pushing forward one of the least studied periods in the history of structure formation. This acquired even more importance in recent times, since it was pointed out by Basu et al. (2024) that our current model of QSO luminosity function evolution yields results that are somewhat in tension with the (still sparse) constraints on the HeEoR.
@@ 5.5 Implications for the relation between sightline effective optical depth and galaxy density
The GaLCC is closely related to the relationship between the total Ly effective optical depth of a sightline and the number density of galaxies around it, that has been recently measured in quasar sightlines (Becker et al., 2018; Kashino et al., 2020; Christenson et al., 2021, 2023; Ishimoto et al., 2022), including some of the most transparent and most opaque LOS known at –. We investigate the implications of the GaLCC and the predictions of thesan regarding this quantity in a companion paper (Garaldi & Bellscheidt, 2024).
@@ 6 Summary and Conclusions
The galaxy–Ly cross-correlation (GaLCC) has been proposed as a novel way to constrain the properties and timing of cosmic reionization. In this paper we have (i) thoroughly investigated the extent to which a different observational limitations prevent us from extracting the intrinsic GaLCC signal, as well as (ii) further clarified the physical origin of the signal and its dependence on galaxy properties and their environment. In order to do so, we have employed the thesan suite of simulations (specifically, the thesan-1 flagship run) to produce synthetic Ly forest lines of sight. Our main results are the following:
@@
We predict the observations by Kashino et al. (2023) to be dominated by stochasticity, and therefore we caution against interpreting them as representative (qualitatively or quantitatively) of the global GaLCC signal. On the contrary, we predict that the latest EIGER results (Kashino et al., 2025), including all 6 sightlines observed, are starting to reliably probe the true GaLCC signal and, therefore, typical IGM conditions. Similarly, the recent results by the ASPIRE program (Kakiichi et al., 2025) span spectral length large enough to beat cosmic variance, but their shallower observations naturally yield a noisier signal.
The galaxy selection method appears to be largely irrelevant for the GaLCC. We do not find differences when the cross-correlation is computed using the galaxies with the largest stellar mass, SFR, gas metallicity (as a proxy for their C IV absorption) and O III flux. This stems from the fact that thesan predicts galaxies at to lie on the galaxy main sequence and the mass-metallicity relation, and the O III flux to be correlated with the galaxy SFR. Hence, galaxies selected to be extreme in one of these quantities are typically extreme also in the others.
@@
The excess transmitted flux at intermediate scales is driven by galaxy overdensities rather than individual bright objects. Therefore, the GaLCC does not trace ionized bubbles produced by individual galaxies, but rather large-scales collective bubbles. This clarifies why Garaldi et al. (2022) found that, simultaneously, (i) even when ionizing photons are allowed to escape only from small () galaxies, the GaLCC signal is not affected, and (ii) the flux excess is larger around massive (, ) galaxies (i.e. when only such galaxies are used to compute the GaLCC). The emerging picture is the following: small galaxies provide the majority of the ionizing photons required to induce an excess Ly transmission, but individual sources are not able to produce an observable signature because of the large ‘structure formation noise’. Therefore, only large groups of such galaxies can produce an observable signature. These groups typically reside in overdense regions, as do highly biased sources like massive galaxies. Hence, using these massive objects to compute the GaLCC effectively selects conglomerates of small galaxies, producing a stronger signal. At the same time, if larger galaxies are not allowed to emit ionizing photons, this does not significantly affect the GaLCC, since the majority of ionising photons is produced by the numerous, nearby smaller galaxies.
The most opaque (transparent) sightlines to Ly photons show larger and closer to the galaxy (smaller and farther from the galaxy) peaks in the GaLCC. This is due to: (i) the fact that the transmitted flux is normalized by its average in the sightline, which is larger in transparent sightlines; and (ii) the overall lower ionising photons density in opaque sightlines renders more difficult to reach the high ionization levels required to allow Ly transmission, and therefore moves the peak transmissivity at smaller distances from the galaxy.
@@
The redshift evolution of the GaLCC peak position is compatible with the evolving density field in a fully-ionized IGM, as expected around the brightest sources of reionization. Quantitatively, the position is approximately times smaller than the mean free path evolution extrapolated from Worseck et al. (2014, which is computed from measurements in the post-reionization IGM).
@@
Overall, our analysis sheds new light on the galaxy–Ly cross correlation. It clarifies further the physical origin and the thoroughly investigate the observational requirements for a faithful determination of this signal, reviewing the ability of current and forthcoming surveys to deliver a faithful measurement. We also provide an extensive discussion of future possible extensions of these type of studies. As such, this paper represents a step further to a deeper understanding of the galaxy–IGM interplay during the first billion years of the Universe, a step direly needed to shed light on the many questions being on a daily basis by observations of the reionizing Universe.
@@ Acknowledgements
EG wishes to express his gratitude towards Koki Kakiichi and Christopher Cain for discussions that inspired part of this work, as well as towards Louise Welsh, Daichi Kashino, Valentina D’Odorico and Matthew Pieri for useful discussions. We are thankful to the community developing and maintaining software packages extensively used in our work, namely: matplotlib (Hunter, 2007), numpy (Walt et al., 2011), scipy (Jones et al., 2001), cmasher (van der Velden, 2020) and CoReCon (Garaldi, 2023).
@@
All simulation data, including snapshots, group and subhalo catalogues, merger trees, and high time cadence Cartesian outputs are publicly available at www.thesan-project.com and thoroughly described in Garaldi et al. (2024), including the additional synthetic spectra developed for this work.
@@
We list here the authors contribution following the CRediT999https://www.elsevier.com/researcher/author/policies-and-guidelines/credit-author-statement system. EG: conceptualization, methodology, software, formal analysis, validation, writing – original draft, writing – review and editing, visualization, supervision, project administration. VB: software, formal analysis, writing – review and editing.
@@ Appendix A Public release of 300 additional sightlines and lightcone-like LOS
| Redshift | - | Redshift of each segment. Notice that this is the redshift the segment would have if the LOS was extracted from a real lightcone (i.e. without piecewise-constant approximation). To know the actual redshift, see the field ‘Snapshot’ |
@@
The 300 additional sightlines have the same data format as the original 150 LOS described in Section 3.10 of Garaldi et al. (2024). They are located at Thesan/Thesan-1/postprocessing/los_zdir. In this directory, there is one file per snapshot containing all LOS. These files are named rays_zdir_NNN.hdf5, where NNN is the 0-left-padded snapshot number (starting from 54).
@@
The lightcone-like LOS are all contained in a single file located at Thesan/Thesan-1/postprocessing/los_lc/rays_lc.hdf5. They share the same format as the fixed-redshift LOS, but contain some additional datasets, reported in Table 2.
Appendix B Impact of peculiar velocities
In Fig. 22 we show the (vanishing) impact of including (fiducial, solid line) or ignoring (dashed line) peculiar velocities when computing the synthetic Ly spectra. As it can be seen clearly, the impact is negligible.
Appendix C Impact of forcing the flux to be positive
@@
In our investigation of the impact of noise on the GaLCC in Sec. 3.3 we do not impose the flux in each pixel to be positive after the addition of noise. This choice stems from the consideration that errors in the quasar continuum estimation can lead to negative flux values. Here we investigate the impact of this choice. Interestingly, we find that forcing the flux to be positive makes the cross-correlation signal much less resilient to increasing noise in the spectrum, simultaneously changing the way it is affected by the noise. We show the impact of forcing the flux in each pixel to be non-negative (essentially by setting all negative flux values to zero) on the cross-correlation in Fig. 23. The difference with Fig. 6 is striking. In this case, the noise does not simply broaden the range of recovered cross-correlation values around the intrinsic one, but instead suppresses the variations (both at small and intermediate scales) and eventually flattens the signal. Therefore, even in the case of moderate S/N (approximately 30%), the recovered signal is far from the intrinsic one. Even more worryingly, the recovered signal appears similar to the intrinsic one at a later time (or, equivalently, larger ionized fraction), potentially preventing the ability to use this probe to pinpoint the completion of the reionization process. The reason of this behaviour change lies in the fact that imposing a positive flux introduces an asymmetry, since now noise can only increase the total flux, unlike in the previous case. Therefore, the noise does not average out when computing the cross-correlation. Instead, it effectively increases the flux regardless of the position of nearby galaxies, hence flattening the cross-correlation.
Appendix D Comparison of effective, median and middle redshift
@@
In Sec. 3.6 we have computed an effective redshift of the GaLCC computed over a broad redshift window by identifying the closest GaLCC computed at fixed redshift (or, equivalently, computed over a narrow redshift range). In Fig. 24 we show how such effective redshift tracks the central redshift of the window used (, top panel) and the median redshift of all galaxies used (, bottom panel), for all combinations of and (the latter shown by the color of the points). In the bottom panel, points are much closer to the one-to-one line (dashed grey line), showing that tracks much better the effective redshift of the cross-correlation computed from the lightcone-like LOS.
@@
Interestingly, appears to almost-fully capture the impact of large redshift windows, as can be seen in the bottom panel of Fig. 24, where all light blue/green points lie close to the one to one line (grey dashed line). This is not true for . At , the scatter around the line increases, indicating that the effective redshift is a worse predictor of than at lower redshift. This is due to the increasing noise in the cross-correlation that makes the identification of an effective redshift more dependent on the noise (arising from the lower number of galaxies fulfilling our selection criteria (i.e. ) decreases at earlier and earlier times. We note that this would be true for any selection criteria based on galaxy mass or brightness.
Appendix E GaLCC evolution fit